OpenStudy (anonymous):

How would you complete 2 itterations of Newtons method using the given initial guess for the following equation: f(x)=2x^2 - 3 x1=1

OpenStudy (da_scienceman):

Write down the Newtons formula, compute the required derivative at initial point and increase the iteration as defined!

OpenStudy (anonymous):

Ok, so \[x _{n+1}=x _{n}-\frac{ f(x _{n}) }{f'(x _{n}) }\] \[f(x)=2x ^{2}-3\] \[f'(x)=4x\] and then ?

OpenStudy (anonymous):

plug in 1 and see what you get

OpenStudy (anonymous):

f(1)=4(1)=4

OpenStudy (anonymous):

\[f(1)=-1, f'(1)=4\] so \[x_2=1-\frac{-1}{4}=\frac{5}{4}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

lather, rinse repeat

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

for what number?

OpenStudy (anonymous):

I got this answer now. Thanks for the help.